### v1

@@ IV.3 Minimisation methods
To minimise the cost function, we use the YOTSE [62] workflow manager, which uses a genetic algorithm. In this, for all parameters, a number of points is drawn from a range between the baseline (Table 2) and their perfect value (except for the coherence time, which is capped at 1 second), i.e., for which $p_{NI}=1$. We initially draw 3 points for server efficiency, 4 for coherence time, 2 for single-qubit fidelity, 3 for entangling gate fidelity and 2 for emission fidelity, meaning the initial population is formed by $3\times 4\times 2\times 3\times 2=144$ sets of parameters. The number of points which are drawn for each parameter is based on the size of the range baseline to perfect for that parameter (e.g., there are more possible values that the entangling gate fidelity can take on than the single-qubit gate fidelity can, as it is further from perfect, we therefore initially draw more points at random from the entangling gate fidelity distribution). From this initial population the lowest-costing 8 ‘parents’ are taken to recombine with a mutation probability of .2 into a new generation. This is repeated over 20 generations. Each point consists of 20000 test rounds for a confidence interval around the error probability of ±$\pm 0.0096$. The minimal costing set of parameters is then fed into a local search algorithm, who decreases the cost of each parameter slightly until the error probability requirement is no longer met. The local search is done with 70000 rounds for a confidence interval of ±$\pm 0.0051$. The outcome of this is a set of parameters that is minimal in the sense that worsening any parameter at that point will result in the requirement not being met.
@@ V Future work
One could modify the cost function (11) to include a constraint on the rate at which the computation or test rounds can be performed. This might change the directions along which the hardware should be improved (i.e., it changes the set of minimal improvements) as parameters such as server efficiency become more important, and conversely the coherence time would become important. This could be done in the same workflow, by changing the cost function and choosing a rate constraint. We have however chosen to not include this, as there is no immediate clear goal in rate, whereas there is a clear goal in error probability (25%).
@@
The constraint on the error probability of 25% we consider in this paper is a theoretical limit. In reality, an error probability of 25% would require an impractically large number of rounds in order to find a desirable $\epsilon$ for security. Instead of setting a constraint on a minimum rate at which the computations can be performed, we might want to find a different metric to target than just the rate and error probability. Having a lower error probability would require one to perform less rounds of the protocol, thereby finishing the total computation faster. An option could be to consider the rate of successfully completed computations, which depends on the error probability as this determines the number of rounds to be performed, as well as on the rate at which these rounds are performed. This is beyond the scope of this paper.
@@
Another thing to consider are requirements for larger graph states. The current analysis considers a universal resource for single-qubit rotations, but a universal resource for any quantum computation (i.e., including two-qubit gates), such as a brickwork state [15], would require more qubits. This will lead to more stringent requirements on the hardware. In principle, the same framework used in this paper could be extended to larger graphs such as the brickwork state, but the current status of the code would make the computation time unreasonably long. The code could be sped up in general by including a framework similar to NetSquid’s entanglement generation Magic [63], which offers simulation speedup through state insertion. An equivalent to this for RSP is currently in development.
@@
The protocol could be optimised by limiting the number of single-qubit gates used in the formation of the graph state, either by considering a full-graph optimisation or by absorbing some of the single-qubit gates into the measurement bases. We also note that, depending on the graph state, not all qubits need to be ‘alive’ in the memory at the same time. The first qubit can be measured before the last qubit is initiated, as long as they are not nearest-neighbours in the graph. In addition, if enough memory qubits are available, the RSP phase could be parallelised by sending ion-entangled photonic qubits from different memory positions successively without waiting for an heralding signal in between.
Data and code availability
The code used to generate, process and plot the data is available at [54], the data is available at [53].
@@ Acknowledgements
We would like to thank Harold Ollivier and Maxime Garnier for useful discussions on rVBQC. We thank Viktor Krutianskii for useful discussions on trapped-ion devices and for sharing experimental data and parameters. We thank David Maier for providing technical support for the code and execution thereof. We thank Kian van der Enden for critical reading of the manuscript. This project (QIA-Phase 1) has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101102140.
@@
JvD led the development and execution of the simulation of the protocol, gathered and analysed the results, and prepared this manuscript. GA contributed to development of hardware models of trapped-ions and prepared Section IV.1 of this manuscript. TzBP, JAS provided feedback and discussions throughout the project. FFdS provided technical support and guidance. TEN helped in the technical understanding of the ion traps and the parameters. All authors revised the manuscript. SW conceived and supervised the project.
@@ Appendix B Alternative route
Multiple sets of parameters can reach the error probability threshold of 25% for a similar cost. In Section III we give a cost-minimised set of parameters (Figure 5) and discuss what what variations in parameters would yield a similar result. The minimised results show less improvement needed in server efficiency and emission fidelity and more in coherence time, entangling and single qubit gate fidelity. How much improvement is needed in these three main objectives varies slightly over different optimisation outcomes. Here, we provide an alternative optimisation outcome compared to the one in the main text to support this observation in Figure 6. In particular, this minimisation puts slightly more emphasis on the single qubit gate fidelity and a little less on the coherence time and entangling gate fidelity compared to the set given in the main text. In addition, the emission fidelity requires more improvement compared to the solution presented in the main text.
@@
Emission fidelity 0.974 0.988
Coherence time [ms] 62 103

### v2

@@ IV.3 Minimisation methods
The genetic algorithm is implemented as follows: for all parameters, a number of points are drawn from a range between the baseline (table 2) and their perfect value (except for the coherence time, which is capped at 1 second), i.e., for which $p_{NI}=1$. We initially draw 3 points for server efficiency, 4 for coherence time, 2 for single-qubit fidelity, 3 for entangling gate fidelity and 2 for emission fidelity, meaning the initial population is formed by $3\times 4\times 2\times 3\times 2=144$ sets of parameters. The number of points which are drawn for each parameter is based on the size of the range baseline to perfect for that parameter (e.g., there are more possible values that the entangling gate fidelity can take on than the single-qubit gate fidelity can, as it is further from perfect; we therefore initially draw more points at random from the entangling gate fidelity distribution). From this initial population, the lowest-costing eight ‘parents’ are taken to recombine with a mutation probability of 0.2 into a new generation. This is repeated over twenty generations. Each point consists of 20000 test rounds for a confidence interval around the error probability of ±$\pm 0.0096$. The set of parameters with minimal cost is then fed into a local search algorithm, who decreases the cost of each parameter slightly until the error probability requirement is no longer met. The local search is done with 70000 rounds for a confidence interval of ±$\pm 0.0051$. The outcome of this is a set of parameters that is minimal in the sense that further parameter adjustments that lower the cost of any parameter at that point will result in the requirement not being met.
Due to the limits of the search space of the parameters, the lowest value for the hardware cost per parameter is one, when the value of this parameter is equal to that of the baseline. Therefore $H_{c}$ is bounded from below by 5 (as 5 parameters are considered). The cost does not have an upper bound, and tends to infinity for any parameter tending to its perfect value. In practice, however, the cost does not extend much past $w_{1}$.
@@ V Future work
One could modify the cost function (10) to include a constraint on the rate at which the computation or test rounds can be performed. This might change the directions along which the hardware should be improved (i.e., it changes the set of minimal improvements) as parameters such as server efficiency become more important, and conversely the coherence time would become less important. This could be done in the same workflow, by changing the cost function and choosing a rate constraint. We have however chosen not to include this, as there is no immediate clear goal in rate, whereas there is a clear goal in error probability (25%).
@@
The constraint on the error probability of 25% we consider in this paper is a theoretical limit. In reality, an error probability of 25% would require an impractically large number of rounds in order to find a desirable $\epsilon$ for security. Instead of setting a constraint on a minimum rate at which the computations can be performed, we might want to find a different metric to target than just the rate and error probability. Having a lower error probability would allow one to perform fewer rounds of the protocol, thereby finishing the total computation faster. An option could be to consider the rate of successfully completed computations, which depends on the error probability, as this determines the number of rounds to be performed, as well as on the rate at which these rounds are performed. This is beyond the scope of this paper.
@@
One could also consider requirements for larger graph states. The current analysis considers a universal resource for single-qubit rotations, but a universal resource for any quantum computation (i.e., including two-qubit gates), such as a brickwork state [16], would require more qubits. This will lead to more stringent requirements on the hardware. In principle, the same framework used in this paper could be extended to larger graphs such as the brickwork state, but the current state of the code would make the computation time quite a bit longer. We estimate that a 10-qubit graph would take about 2.5 times as long as the 5-qubit graph, the method for this estimation is described in C. This means that the full optimisation procedure (i.e., 20 generations of the genetic algorithm) is estimated to take about 20000 core-hours to complete for a 10-qubit version. The code could be sped up by including a framework similar to NetSquid’s entanglement generation Magic [75], which offers simulation speedup through state insertion. An equivalent to this for RSP is currently in development.
@@
The protocol could be optimised by limiting the number of single-qubit gates used in the formation of the graph state, either by considering a full-graph optimisation or by absorbing some of the single-qubit gates into the measurement bases. We also note that, depending on the graph state, not all qubits need to be ‘alive’ in the memory at the same time. The first qubit can be measured before the last qubit is initiated, as long as they are not nearest-neighbours in the graph. In addition, if additional memory qubits are available, the RSP phase could be parallelised by sending ion-entangled photonic qubits from different memory positions successively without waiting for a heralding signal in between.
VI Data and code availability
The code used to generate, process and plot the data is available at [76], and the data is available at [63].
@@ Acknowledgements
We thank Harold Ollivier and Maxime Garnier for useful discussions on rVBQC. We thank Viktor Krutianskii for useful discussions on trapped-ion devices and for sharing experimental data and parameters. We thank David Maier for providing technical support for the code and execution thereof. We thank Kian van der Enden for critical reading of the manuscript. This project (QIA-Phase 1) has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101102140.
@@
JvD led the development and execution of the simulation of the protocol, gathered and analysed the results, and prepared this manuscript. GA contributed to development of hardware models of trapped ions and prepared section IV.1 of this manuscript. TzBP and JAS provided feedback and discussions throughout the project. FFdS provided technical support and guidance. TEN helped in the technical understanding of the ion traps and the parameters. All authors revised the manuscript. SW conceived and supervised the project.
@@ Appendix B Alternative route
Multiple sets of parameters can reach the error probability threshold of 25% for a similar cost. In Section III we give a cost-minimised set of parameters (Figure 5) and discuss what what variations in parameters would yield a similar result. The minimised results show less improvement needed in server efficiency and emission fidelity and more in coherence time, entangling and single qubit gate fidelity. How much improvement is needed in these three main objectives varies slightly over different optimisation outcomes. Here, we provide an alternative optimisation outcome compared to the one in the main text to support this observation. In particular, this minimisation puts slightly more emphasis on the single qubit gate fidelity and a little less on the coherence time and entangling gate fidelity compared to the set given in the main text. In addition, the emission fidelity requires more improvement compared to the solution presented in the main text.
@@
Ion-photon entanglement fidelity 0.974 0.988
@@
Appendix C Estimation of runtime for larger graph states
The simulation currently simulates every attempt on remote state preparation; also the failed ones. When the size of the graph gets larger (i.e., it contains more qubits) we will both need to have success more often and, in particular, they all need to happen within the cutoff time, which is half of the coherence time. This will take longer to simulate. The full optimisation procedure (i.e., 20 generations of the genetic algorithm) has not been run for larger graph sizes. To make an estimate of how long this would take, we can compare how long the 5-qubit optimisation procedure took to the fractional runtime of larger graphs at the baseline. When computation B takes twice as long to perform as computation A, the fractional runtime of B compared to A is 2. The optimisation for the 5-qubit graph took roughly 8000 core-hours, a computation with a fractional runtime of 2 compared to the 5-qubit graph would then take roughly 16000 core-hours.
To make an estimate of the fractional runtime, we ran the baseline parameters for multiple graph sizes for 5000 test rounds, recorded the time it took to finish and compared this to the time it took to finish the same calculation with a 5-qubit graph. The result of this can be found in figure 7. The fractional runtime seems to increase roughly linearly with increasing number of qubits in the graph, so a linear fit was included to possibly extend this logic to slightly larger graphs. The linear fit results in the relation $y=0.281x-0.376$ with $x$ the number of qubits in the graph and $y$ the fractional runtime. However, we do not expect this relation to be linear for all graph sizes as we expect that the probability of all qubits being remotely prepared withing the cutoff time will go to zero more quickly for larger graphs. Additionally, this is a rough estimate as it is only fully applicable to the hardware parameters at baseline, whereas the optimisation procedure will have different parameter sets, though we hope it provides a useful estimate nonetheless.
Figure 7: Estimation of the fractional runtime (y-axis) for calculations including a certain number of qubits in the graph (x-axis) compared to the calculation involving a 5-qubit graph. Plot includes standard error of the mean for 5000 points per graph size, error bars smaller than markers (blue). Also plotted with linear fit (orange). This can be used to estimate how long the full optimisation procedure will take.